# OptiPMB: Enhancing 3D Multi-Object Tracking with Optimized Poisson Multi-Bernoulli Filtering

### v1

@@ IV-G Adaptive Designs
IV-G1 Adaptive Detection Probability
@@
IV-G2 Hybrid Adaptive Birth Model
@@
$\mathbf{m}_{k}^{0,m}=\begin{bmatrix}z_{x,k}^{m}\\ z_{y,k}^{m}\\ \sqrt{(z_{vx,k}^{m})^{2}+(z_{vy,k}^{m})^{2}}\\ \arctan(z_{vy,k}^{m},z_{vx,k}^{m})\\ 0\\ 0\end{bmatrix},\ \ P_{k}^{0,m}=P^{0}(z_{c,k}^{m}).$ (39)
@@ IV-H Optimizations
IV-H1 Redundant PPP Pruning
Although the HABM proposed in OptiPMB can reduce false negative errors by introducing the measurement-driven newborn PPP intensity $\lambda_{k}^{\mathrm{b}}(\mathbf{x})$, the number of Poisson components within $\lambda_{k}^{\mathrm{b}}(\mathbf{x})$ increases over time. To improve computational efficiency and reduce false track initialization, removing redundant Poisson components during the PPP pruning step is necessary. While the previous PMBM-based 3D MOT methods [19, 20] did not specify a dedicated pruning strategy, the original PMBM filter [27] only prunes the Poisson components with weights below a predefined threshold. However, this approach is not well-suited for measurement-driven PPP birth modeling. Since the PPP removal solely depends on one pruning threshold, new Poisson components with low initial weights may be pruned before they can correctly initialize a new track, while Poisson components that have already generated detected potential objects may persist unnecessarily, leading to false track initiations. The GNN-PMB tracker [21] utilizes a full measurement-driven approach for generating $\lambda_{k}^{\mathrm{b}}(\mathbf{x})$ but removes all Poisson components at the pruning step. This aggressive pruning strategy may result in track initialization failures.
@@ IV-H1 Redundant PPP Pruning
IV-H2 Optimized Track Extraction
@@
TABLE I: Finetuned OptiPMB parameters.
Parameters nuScenes Category KITTI
Bic. Bus Car Mot. Ped. Tra. Tru. Car
Score filter threshold $\eta_{\mathrm{sf}}$ 0.15 0 0.1 0.16 0.2 0.1 0 0.6
NMS IoU threshold $\eta_{\mathrm{iou}}$ 0.1 0.1 0.1 0.1 0.1 0.1 0.1 0.1
Survival probability $p_{\mathrm{s}}$ 0.99 0.99 0.99 0.99 0.99 0.99 0.99 0.999
Gating distance threshold $\eta_{\mathrm{dist}}$ 3 10 10 4 3 10 10 4
Expected number of points $\mathrm{PTS}_{0}$ 10 10 10 10 10 10 10 20
Minimal scaling factor $s_{\mathrm{d}}$ 0.5 0.5 0.5 0.5 0.5 0.5 0.5 0.5
HABM score threshold $\eta_{\mathrm{score}}$ 0.17 0.3 0.25 0.18 0.2 0.15 0.15 0.85
Adaptive birth rate $\mu^{\mathrm{ab}}$ 2 2 2 2 2 2 2 2
Undetected object birth rate $\mu^{\mathrm{b0}}$ 1 5 2 1 1 2 2 1
Clutter rate $\mu^{\mathrm{c}}$ 0.5 0.2 1 0.5 0.5 0.5 1 5
PPP max survival time steps $\eta_{\mathrm{step}}$ 3 3 3 2 2 2 2 4
Track extraction threshold 1 $\eta_{\mathrm{ext1}}$ 0.7 0.7 0.7 0.7 0.7 0.7 0.5 0.95
Track extraction threshold 2 $\eta_{\mathrm{ext2}}$ 0.95 0.7 0.8 0.95 0.8 0.8 0.9 0.98
Misdetection counter limit $\eta_{\mathrm{cnt}}$ 3 2 2 2 2 2 2 3
[…]
Overall Bic. Bus Car Mot. Ped. Tra. Tru.
RFS-M3 (ICRA 2021) [20] CenterPoint L 0.619 N/A N/A N/A N/A N/A N/A N/A 0.752 90872 27168 16728 1525
OGR3MOT (RA-L 2022) [1] CenterPoint L 0.656 0.380 0.711 0.816 0.640 0.787 0.671 0.590 0.620 95264 24013 17877 288
SimpleTrack (ECCVW 2022) [10] CenterPoint L 0.668 0.407 0.715 0.823 0.674 0.796 0.673 0.587 0.550 95539 23451 17514 575
GNN-PMB (T-IV 2023) [21] CenterPoint L 0.678 0.337 0.744 0.839 0.657 0.804 0.715 0.647 0.560 97274 21521 17071 770
3DMOTFormer (ICCV 2023) [4] CenterPoint L 0.682 0.374 0.749 0.821 0.705 0.807 0.696 0.626 0.496 95790 23337 18322 438
NEBP (T-SP 2023) [49] CenterPoint L 0.683 0.447 0.708 0.835 0.698 0.802 0.690 0.598 0.624 97367 21971 16773 227
ShaSTA (RA-L 2024) [3] CenterPoint L 0.696 0.410 0.733 0.838 0.727 0.810 0.704 0.650 0.540 97799 21293 16746 473
OptiPMB (Ours) CenterPoint L 0.713 0.414 0.748 0.860 0.751 0.836 0.727 0.658 0.541 98550 20758 16566 257
Poly-MOT (IROS 2023) [13] LargeKernel3D C+L 0.754 0.582 0.786 0.865 0.810 0.820 0.751 0.662 0.422 101317 17956 19673 292
Fast-Poly (RA-L 2024) [14] LargeKernel3D C+L 0.758 0.573 0.767 0.863 0.826 0.852 0.768 0.656 0.479 100824 18415 17098 326
${}^{\text{\textdagger}}$ (Ours) LargeKernel3D C+L 0.767 0.600 0.774 0.870 0.827 0.848 0.775 0.676 0.460 102899 16411 18796 255
@@ V-A Dataset and Implementation Details
TABLE III: Online 3D MOT performance of OptiPMB and other state-of-the-art trackers on the nuScenes validation set. The CenterPoint [69] detector is employed for all trackers for a fair comparison. Bold and underline denote the first and second best results. Blue indicates the two primary tracking accuracy metrics, HOTA and AMOTA.
[…]
${}^{\text{\textdagger}}$ (IEEE T-IV 2023) [21] Bic. 17.88 4.82 66.54 0.513 0.602 1135 855 220 3
Bus 46.83 30.16 72.97 0.854 0.546 1789 312 155 11
Car 57.53 43.22 77.37 0.849 0.387 49182 8791 6140 344
Mot. 21.38 6.91 66.22 0.723 0.475 1494 477 247 6
Ped. 46.49 29.86 72.48 0.807 0.357 21149 4008 3669 266
Tra. 23.90 9.03 64.22 0.506 0.973 1378 1045 412 2
Tru. 36.13 17.38 75.36 0.695 0.581 7007 2625 1519 18
Overall 46.03 28.15 75.45 0.707 0.560 83134 18113 12362 650
${}^{\text{\textdagger}}$ (IEEE RA-L 2024) [3] Bic. 13.62 2.58 72.12 0.588 0.507 1187 804 143 2
Bus 45.17 26.85 76.16 0.856 0.527 1777 333 142 2
Car 57.05 41.48 78.67 0.856 0.337 49200 8986 6067 131
Mot. 21.56 6.83 68.10 0.745 0.521 1457 513 83 7
Ped. 46.99 29.48 74.95 0.814 0.378 20945 4258 3378 220
Tra. 23.69 8.63 65.84 0.531 0.970 1284 1139 238 2
Tru. 36.24 17.17 76.70 0.703 0.572 7222 2421 1671 7
Overall 43.87 24.98 77.18 0.728 0.544 83072 18454 11722 371
${}^{\text{\textdagger}}$ (IEEE RA-L 2024) [14] Bic. 16.36 3.61 74.30 0.572 0.553 1218 775 189 0
Bus 46.30 27.97 76.90 0.855 0.522 1787 316 172 9
Car 62.63 49.48 79.56 0.860 0.339 50443 7654 8057 220
Mot. 27.44 9.95 75.83 0.792 0.470 1559 415 178 3
Ped. 57.59 43.05 77.12 0.835 0.355 21916 3332 4070 175
Tra. 24.69 9.29 66.40 0.529 0.939 1440 984 371 1
Tru. 32.29 13.40 78.00 0.714 0.533 7220 2424 1676 6
Overall 48.52 30.04 78.55 0.737 0.530 85583 15900 14713 414
OptiPMB (Ours) Bic. 17.50 4.20 73.07 0.575 0.514 1221 771 208 1
Bus 55.91 40.69 77.27 0.857 0.538 1785 327 131 0
Car 62.69 49.10 80.47 0.870 0.359 50396 7864 6850 57
Mot. 32.52 14.06 75.39 0.792 0.470 1497 477 112 3
Ped. 61.73 49.41 77.24 0.838 0.357 21365 3892 3401 166
Tra. 27.12 10.59 70.70 0.534 0.949 1488 937 399 0
Tru. 38.77 19.05 79.21 0.718 0.539 7211 2439 1576 1
Overall 52.06 34.31 79.27 0.741 0.532 84963 16707 12677 227
  • Evaluated using the authors’ parameters. The tracking performance is identical to that reported in the original manuscripts [21, 3, 14].
@@ V-B Comparison with State-of-the-Arts
As shown in Table II, our proposed OptiPMB demonstrates state-of-the-art online 3D MOT performance on the nuScenes test set. Specifically, it achieves overall AMOTA scores of 0.767 using the LargeKernel3D [71] detector and 0.713 using the CenterPoint [69] detector. With the same detectors, OptiPMB outperforms the previous learning-assisted [1, 49], learning-based [4, 3] and model-based [20, 10, 21, 13, 14] methods in terms of the primary tracking accuracy metric, AMOTA. Our method also demonstrates strong performance in secondary metrics, including TP, FN, FP, and IDS, especially comparing with the previous model-based RFS trackers, the RFS-M3 [20] and GNN-PMB [21]. Our method is fully open-sourced, with code and parameters available, to demonstrate the potential of RFS-based trackers and serve as a new baseline for model-based 3D MOT.
@@
TABLE IV: Online 3D MOT performance of OptiPMB and other advanced trackers on the KITTI car validation set, following the evaluation protocol in [8]. Trackers using PointRCNN [72] and CasA [70] detectors are distinguished for fair comparison, with different modalities (C for camera, L for LiDAR) indicated. Bold and underline denote the first and second best results among trackers using the same detector.
@@
OptiPMB (Ours) PointRCNN L 93.77 47.56 76.77 87.99 77.11 0
${}^{\text{\textdagger}}$ (arXiv 2024) [15] CasA L 91.69 47.29 81.88 90.52 84.01 0
${}^{\text{\textdagger}}$ (T-GRS 2022) [70] CasA L 92.14 44.94 82.60 89.45 84.34 2
OptiPMB (Ours) CasA L 94.07 47.47 81.77 91.00 81.20 0
  • Evaluated using online tracking configurations and the authors’ parameters. The tracking performance is identical to that reported by the authors.
TABLE V: Online 2D MOT performance of our proposed OptiPMB and other advanced trackers on the KITTI car test dataset. Trackers using PointRCNN [72] and CasA [70] detectors are distinguished for fair comparison, with different modalities (C for camera, L for LiDAR) indicated. Bold and underline denote the first and second best results among trackers using the same detector. Blue indicates the primary tracking accuracy metric, HOTA.
@@
AB3DMOT (IROS 2020) [8] PointRCNN L 69.99 71.13 69.33 83.61 85.23 29849 4543 113
Feng et al. (T-IV 2024) [34] PointRCNN C+L 74.81 N/A N/A 84.82 85.17 N/A N/A N/A
DeepFusionMOT (RA-L 2022) [12] PointRCNN C+L 75.46 71.54 80.05 84.63 85.02 29791 4601 84
StrongFusionMOT (S-J 2023) [35] PointRCNN C+L 75.65 72.08 79.84 85.53 85.07 29734 4658 58
OptiPMB (Ours) PointRCNN L 74.04 67.46 81.69 77.47 84.79 31506 2886 61
UG3DMOT (SP 2024) [29] CasA L 78.60 76.01 82.28 87.98 86.56 31399 2992 30
MMF-JDT (RA-L 2025) [39] CasA C+L 79.52 75.83 84.01 88.06 86.24 N/A N/A 37
OptiPMB (Ours) CasA L 80.26 76.93 84.36 88.92 86.66 32353 2039 43
V-B2 KITTI Dataset
Since the official KITTI object tracking evaluation only includes 2D MOT metrics, we compare the online 3D MOT performance of OptiPMB and other advanced trackers on the car category of the KITTI validation dataset, following the evaluation protocol proposed in [8]. This protocol designates sequences 1, 6, 8, 10, 12, 13, 14, 15, 16, 18, and 19 in the original KITTI training set as the validation set. As shown in Table IV, OptiPMB achieves superior tracking accuracy on the KITTI validation set, outperforming other state-of-the-art trackers that use the same PointRCNN [72] and CasA [70] detectors across sAMOTA, AMOTA, and MOTA metrics. Here, the online 3D MOT performance of MCTrack [15] and CasTrack [70] is evaluated using their online tracking configurations and finetuned parameters provided by the authors. For reference, the 2D MOT performance for the compared trackers on the KITTI test set is presented in Table V. With the PointRCNN [72] detection results, OptiPMB achieves significantly higher tracking accuracy than the baseline AB3DMOT [8] tracker (+4.05% HOTA) while generating the fewest false positives among the compared methods. Since OptiPMB relies solely on 3D detection results, camera-LiDAR fusion-based trackers [34, 12, 35] attain higher HOTA than it, as expected. However, when utilizing higher-quality CasA [70] detection results, OptiPMB surpasses even the recently proposed camera-LiDAR fusion-based MMF-JDT [39] tracker in 2D MOT performance.
@@ V-C Ablation Study on OptiPMB
We perform ablation experiments on the nuScenes validation set to evaluate the impact of various components proposed in OptiPMB, as shown in Table VI. The tracker variants share the same parameters if not specifically mentioned.
@@
Remove ${}^{\text{\textdagger}}$ 0.732 0.563 17595 11391 264
Remove ${}^{\text{\textdagger}}$ 0.709 0.614 19420 12033 279
Remove ${}^{\text{\textdagger}}$ 0.739 0.534 16516 12911 233
Remove ${}^{\text{\textdagger}}$ 0.722 0.529 17678 15046 163
OptiPMB (Ours) 0.741 0.532 16707 12677 227
  • ADP: adaptive detection probability. HABM: hybrid adaptive birth model. RPP: redundant PPP pruning. OTE: optimized track extraction.
V-C1 Effectiveness of the Proposed Adaptive Designs
After replacing the adaptive detection probability $p_{\mathrm{d}}(\mathbf{x}_{+})$ with the fixed parameter $p_{\mathrm{d0}}$, the tracker produces more FN and IDS errors due to occlusions, leading to lower overall tracking accuracy (-0.9% AMOTA). The removal of HABM significantly impacts the initialization of new tracks (+2713 FN), reducing the tracking accuracy to a level comparable with GNN-PMB (-3.2% AMOTA). This suggests that the proposed HABM is a primary contributor to the effectiveness of OptiPMB.
V-C2 Effectiveness of the Proposed Optimizations
As shown in Table VI, the redundant PPP pruning module contributes less significantly to tracking performance when the other three modules are present. However, its absence leads to a decrease in tracking accuracy (-0.2% AMOTA) and an increase in false positives (+234 FP). Without the optimized track extraction module, the tracker extracts object tracks with existence probabilities exceeding an average threshold, $(\eta_{\mathrm{ext}1}+\eta_{\mathrm{ext}2})/2$. This modification slightly improves track ID maintenance but worsens both track initialization and termination performance (+971 FN, +2369 FP), resulting in a lower overall tracking accuracy (-1.9% AMOTA).
@@ V-C2 Effectiveness of the Proposed Optimizations
V-D Parameter Sensitivity and Runtime
The performance of model-based 3D MOT methods often relies on careful selection of hyperparameters. Therefore, we analyzed the impact of key parameters in OptiPMB on the car category in the nuScenes validation set. As shown in Fig. 5, the survival probability ($p_{\mathrm{s}}$), clutter rate ($\mu^{\mathrm{c}}$), and misdetection counter threshold ($\eta_{\mathrm{cnt}}$) significantly influence both AMOTA and HOTA scores. To achieve better track continuity and recall performance without sacrificing tracking accuracy, we suggest setting these parameters to $p_{\mathrm{s}}=0.99$, $\mu^{\mathrm{c}}=1$, and $\eta_{\mathrm{cnt}}=2$. OptiPMB is less sensitive to other hyperparameters, demonstrating the robustness of our proposed adaptive designs. On our test platform with the Ubuntu 22.04 operating system, an AMD 7950X CPU, and 64GB of RAM, the average processed frames per second (FPS) for the Python implementation of OptiPMB on the nuScenes and KITTI validation sets are 11.3 FPS and 122.1 FPS, respectively.

### v2

@@ IV-G Adaptive Designs
IV-G1 Adaptive Detection Probability (ADP)
@@
IV-G2 Hybrid Adaptive Birth Model (HABM)
@@
$\begin{cases}\mathbf{m}_{k}^{0,m}=[z_{x,k}^{m},z_{y,k}^{m},\sqrt{(z_{vx,k}^{m})^{2}+(z_{vy,k}^{m})^{2}},\\ \qquad\qquad\qquad\qquad\arctan(z_{vy,k}^{m},z_{vx,k}^{m}),0,0]^{\mathrm{T}}\\ P_{k}^{0,m}=P^{0}(z_{c,k}^{m}).\end{cases}$ (39)
@@ IV-H Optimizations
IV-H1 Redundant PPP Pruning (RPP)
Although the HABM proposed in OptiPMB can reduce false negative errors by introducing the measurement-driven newborn PPP intensity $\lambda_{k}^{\mathrm{b}}(\mathbf{x})$, the number of Poisson components within $\lambda_{k}^{\mathrm{u}}(\mathbf{x})$ increases over time. To improve computational efficiency and reduce false track initialization, removing redundant Poisson components during the PPP pruning step is necessary. While the previous PMBM-based 3D MOT methods [20, 21] did not specify a dedicated pruning strategy, the original PMBM filter [60] only prunes the Poisson components with weights below a predefined threshold. However, this approach is not well-suited for measurement-driven PPP birth modeling. Since the PPP removal solely depends on one pruning threshold, new Poisson components with low initial weights may be pruned before they can correctly initialize a new track, while Poisson components that have already generated detected potential objects may persist unnecessarily, leading to false track initiations. The GNN-PMB tracker [22] utilizes a full measurement-driven approach for generating $\lambda_{k}^{\mathrm{b}}(\mathbf{x})$ but removes all Poisson components at the pruning step. This aggressive pruning strategy may result in track initialization failures.
@@ IV-H1 Redundant PPP Pruning (RPP)
IV-H2 Optimized Track Extraction (OTE)
@@ V-A Dataset and Implementation Details
TABLE I: Finetuned OptiPMB parameters.
Parameters nuScenes Category KITTI
Bic. Bus Car Mot. Ped. Tra. Tru. Car
Score filter threshold $\eta_{\mathrm{sf}}$ 0.15 0 0.1 0.16 0.2 0.1 0 0.3
NMS IoU threshold $\eta_{\mathrm{iou}}$ 0.1 0.1 0.1 0.1 0.1 0.1 0.1 0.1
Survival probability $p_{\mathrm{s}}$ 0.99 0.99 0.99 0.99 0.99 0.99 0.99 1
Gating distance threshold $\eta_{\mathrm{dist}}$ 3 10 10 4 3 10 10 4
Expected number of points $\mathrm{PTS}_{0}$ 10 10 10 10 10 10 10 10
Minimal scaling factor $s_{\mathrm{d}}$ 0.5 0.5 0.5 0.5 0.5 0.5 0.5 0.7
HABM score threshold $\eta_{\mathrm{score}}$ 0.17 0.3 0.25 0.18 0.2 0.15 0.15 0.65
Adaptive birth rate $\mu^{\mathrm{ab}}$ 2 2 2 2 2 2 2 0.2
Undetected object birth rate $\mu^{\mathrm{b0}}$ 1 5 2 1 1 2 2 0.5
Clutter rate $\mu^{\mathrm{c}}$ 0.5 0.2 1 0.5 0.5 0.5 1 10
PPP max survival time steps $\eta_{\mathrm{step}}$ 3 3 3 2 2 2 2 1
Track extraction threshold 1 $\eta_{\mathrm{ext1}}$ 0.7 0.7 0.7 0.7 0.7 0.7 0.5 0.85
Track extraction threshold 2 $\eta_{\mathrm{ext2}}$ 0.95 0.7 0.8 0.95 0.8 0.8 0.9 0.9
Misdetection counter limit $\eta_{\mathrm{cnt}}$ 3 2 2 2 2 2 2 5
@@ V-B Comparison with State-of-the-Arts
As shown in Table II, our proposed OptiPMB demonstrates state-of-the-art online 3D MOT performance on the nuScenes test set. Specifically, it achieves overall AMOTA scores of 0.767 using the LargeKernel3D [69] detector and 0.713 using the CenterPoint [67] detector. With the same detectors, OptiPMB outperforms the previous learning-assisted [1, 44], learning-based [4, 3] and model-based [21, 10, 22, 13, 14] methods in terms of the primary tracking accuracy metric, AMOTA. Our method also demonstrates strong performance in secondary metrics, including TP, FN, FP, and IDS, especially comparing with the previous model-based RFS trackers, the RFS-M3 [21] and GNN-PMB [22]. Our method is fully open-sourced, with code and parameters available [^3] Code is available at: https://github.com/dinggh0817/OptiPMB, to demonstrate the potential of RFS-based trackers and serve as a new baseline for model-based 3D MOT.
@@
[…]
TABLE III: Online 3D MOT performance of OptiPMB and other state-of-the-art trackers on the nuScenes validation set. The CenterPoint [67] detector is employed for all trackers for a fair comparison. Bold and underline denote the first and second best results within each category. Blue indicates the two primary tracking accuracy metrics, HOTA and AMOTA.
[…]
  • Evaluated using the authors’ code and parameters. The tracking performance is identical to that reported in the original manuscripts.
@@
V-B2 KITTI Dataset
Since the official KITTI object tracking evaluation only includes 2D MOT metrics, we compare the online 3D MOT performance of OptiPMB and other advanced trackers on the car category of the KITTI validation dataset, following the evaluation protocol proposed in [8]. This protocol designates sequences 1, 6, 8, 10, 12, 13, 14, 15, 16, 18, and 19 in the original KITTI training set as the validation set. As shown in Table IV, OptiPMB achieves superior tracking accuracy on the KITTI validation set, outperforming other state-of-the-art trackers that use the same PointRCNN [70], PointGNN [71], and CasA [68] detectors in sAMOTA scores. Here, the online 3D MOT results of CasTrack [68], MCTrack [15], and RobMOT [72] are evaluated using their online tracking configurations and finetuned parameters provided by the authors. For reference, the 2D MOT performance for the compared trackers on the KITTI test set is presented in Table V. With the PointRCNN [70] and PointGNN [71] detection results, OptiPMB achieves significantly higher tracking accuracy than the baseline LiDAR-only trackers [8, 50] while generating the fewest false positives and ID switch errors among the compared methods. Although OptiPMB relies solely on LiDAR detection results, its HOTA score is still comparable with camera-LiDAR fusion-based trackers [29, 12, 30, 2]. Moreover, when utilizing higher-quality CasA [68] detection results, OptiPMB surpasses even the recently proposed RobMOT [72] and MMF-JDT [76] trackers in HOTA, DetA, MOTA, TP, and FP scores, demonstrating a strong tracking performance.
TABLE IV: Online 3D MOT performance of OptiPMB and other advanced trackers on the KITTI car validation set. PointRCNN [70], Point-GNN [71], and CasA [68] detectors are distinguished for fair comparison, with different modalities (C for camera, L for LiDAR) indicated. Bold and underline denote the first and second best results among trackers using the same detector.
@@ V-B2 KITTI Dataset
OptiPMB (Ours) PointRCNN L 93.78 48.40 77.30 87.53 77.39 0
PolarMOT (ECCV 2022) [50] Point-GNN L 94.32 N/A N/A 93.93 N/A 31
CAMO-MOT (T-ITS 2023) [2] Point-GNN C+L 95.20 48.04 81.48 N/A N/A N/A
OptiPMB (Ours) Point-GNN L 96.29 48.87 79.39 93.48 78.75 0
RobMOT $^{\text{\textdagger}}$ (T-ITS 2025) [72] CasA L 91.16 46.95 82.06 89.53 83.57 0
@@
[…]
OptiPMB (Ours) CasA L 93.40 47.19 81.47 91.00 81.03 0
  • Evaluated using the authors’ official code, parameters, and online tracking configurations.
TABLE V: Online 2D MOT performance of our proposed OptiPMB and other advanced trackers on the KITTI car test dataset. Trackers using PointRCNN [70], Point-GNN [71], and CasA [68] detectors are distinguished for fair comparison, with different modalities (C for camera, L for LiDAR) indicated. Bold and underline denote the first and second best results among trackers using the same detector. Blue indicates the primary tracking accuracy metric, HOTA.
[…]
@@ V-C Ablation Study on OptiPMB
We conduct ablation experiments on the nuScenes validation set to assess the influence of the proposed components in OptiPMB, as summarized in Table VI. In the baseline tracker, the adaptive detection probability $p_{\mathrm{d}}(\mathbf{x}_{+})$ is replaced with the fixed parameter $p_{\mathrm{d0}}$, and object tracks with existence probabilities exceeding the threshold $\eta_{\mathrm{ext}1}$ are extracted. In addition, the HABM and OTE modules are replaced by the object birth model and the PPP pruning strategy employed in the GNN-PMB [22] tracker. Except for the detection probability and existence threshold settings described above, all tracker variants share identical parameter configurations. For clarity of comparison, trackers equipped with different numbers of modules are grouped together and ranked within each group according to their AMOTA scores. To analyze the runtime impact of different modules, we also report the average processed frames per second (FPS) of the Python implementation of each tracker variant on our test platform, running Ubuntu 22.04 with an AMD 7950X CPU and 64 GB of RAM. Consistent with common practice in the TBD literature, these FPS values are measured without accounting for detector latency.
As shown in Table VI, HABM is the primary contributor to the effectiveness of OptiPMB, as trackers equipped with HABM achieve the highest AMOTA scores within each group. Incorporating this birth model not only substantially boosts the baseline tracking performance (+3.0% AMOTA), but also improves both track ID maintenance and runtime efficiency (–36 IDS, +2.6 FPS) by generating more reliable newborn objects and suppressing false potential objects. However, the other three modules exhibit more complex effects and cannot effectively improve tracking accuracy in isolation. Specifically, adding ADP reduces ID switch errors by enhancing the maintenance of occluded tracks, but it also introduces additional false tracks, which can reduce tracking accuracy. The RPP module reduces false negatives by retaining PPP components for a longer duration, but at the cost of increased tracking latency due to the larger number of potential objects. The OTE module improves accuracy and FPS by reducing false tracks, yet simultaneously increases ID switch errors. Nevertheless, when combined with HABM, the adverse effects of these modules are mitigated, leading to the improvement in overall tracking performance. Consequently, our OptiPMB tracker equipped with all four modules achieves the highest tracking accuracy (+4.5% AMOTA) among all tracker variants, demonstrating the effectiveness of their joint design.
@@
0.696 0.622 20421 12362 250 17.5
0.688 0.609 21458 13755 178 15.4
0.692 0.614 19664 12683 325 4.7
0.701 0.627 20132 11020 334 19.0
0.726 0.551 16061 14403 214 20.1
0.688 0.597 20985 13820 203 4.3
0.697 0.632 19941 11082 414 4.7
0.705 0.621 20103 11301 275 14.6
0.721 0.535 17766 14975 163 13.2
0.727 0.551 17432 13097 206 17.2
0.731 0.569 17607 11385 264 19.7
0.708 0.615 19700 11214 414 4.3
0.722 0.529 17678 15046 163 13.3
0.732 0.563 17595 11391 264 15.6
0.739 0.534 16516 12911 233 13.4
0.741 0.532 16707 12677 227 11.3
@@
V-D Parameter Sensitivity
The performance of model-based 3D MOT methods often relies on careful selection of hyperparameters. Therefore, we analyzed the impact of key parameters in OptiPMB on the car category in the nuScenes validation set. As shown in Fig. 5, the survival probability ($p_{\mathrm{s}}$), clutter rate ($\mu^{\mathrm{c}}$), and misdetection counter threshold ($\eta_{\mathrm{cnt}}$) significantly influence both AMOTA and HOTA scores. To achieve better track continuity and recall performance without sacrificing tracking accuracy, we suggest setting these parameters to $p_{\mathrm{s}}=0.99$, $\mu^{\mathrm{c}}=1$, and $\eta_{\mathrm{cnt}}=2$. OptiPMB is less sensitive to other hyperparameters, demonstrating the robustness of our proposed adaptive designs.